# Fast Inference Using Automatic Differentiation and Neural Transport in Astroparticle Physics

### v1

@@ Results
To compare the performance of NUTS and NS, both with and without NeuTra, we use the metrics of wall time per effective sample size (ESS), and likelihood or gradient evaluation per ESS. For NUTS, as ESS can be different for each parameter, the minimum ESS is used, following hoffman2014no . The calculation of ESS for MCMC and NS is explained in Appendix D. We note that the ESS per evaluation metric is not directly comparable between NS and NUTS because gradient evaluations are required for NUTS and not for NS. We also include the number of divergences encountered per ESS for NUTS runs, as these indicate regions of high curvature that often cannot be reliably sampled by MCMC Betancourt:2017ebh .
We show the marginalized posteriors for $\mu_{1}$ and $\mathcal{C}_{1}$ in Fig. 2, produced using corner corner . We see that the $68\%$ and $95\%$ highest posterior density region contours, as well as the probability densities of these parameters, all show a high degree of overlap. Our results also agree with the ground truth. The full corner plot for all parameters is shown in Appendix E, and they exhibit the same behaviors.
@@
We also find that the evidence computed using nested sampling, shown in Appendix F, matches the ground truth value of $-11.14$. The mean ELBO from the last 100 epochs of neural transport training, $-11.21$, is also close to the ground truth value. This demonstrates that NeuTra can be applied to NS to accelerate Bayesian evidence computation and model comparison.
3.2 Application to Neutrino Physics
Neutrino Physics Model
@@ 3.2 Application to Neutrino Physics
For our physics application, we construct our likelihood using the NR result from XENON1T XENON:2020gfr and the ER measurement from PandaX-4T Lu:2024ilt . These measurements probe complementary parts of the NSI parameter space, allowing us to make stronger inferences compared to using either result alone.
@@
For the XENON1T NR measurement, the data contains both solar neutrino events and background events that are unrelated to our signal of interest. As our likelihood model, we use a Poisson distribution with expectation value $\lambda\equiv\lambda_{\mathrm{NR}}+\lambda_{\mathrm{bkg}}$. Here, $\lambda_{\text{NR}}$ is the expected number of NR events, calculated using SNuDD via Eq. 6, and $\lambda_{\mathrm{bkg}}$ is the expected number of background events. As given in XENON:2020gfr , we use $\lambda_{\mathrm{bkg}}=5.38$ and take the number of observed NR events to be $N_{\mathrm{NR}}=6$.
@@
For the PandaX-4T ER result, we use their inferred result on the number of solar neutrinos arriving at the detector per second per unit area, which is proportional to the number of ER events. As in the NR case, we use SNuDD to compute the number of expected ER events for a given set of NSI parameters. We then compare this number to the one calculated in the absence of novel phenomena using the ratio between the two values, $r_{\mathrm{ER}}$. We model the distribution of this ratio as a normal distribution truncated at zero at the lower end with mean and standard deviation both given by $1.72$. The truncation is to indicate that this physical quantity is strictly positive. The value $1.72$ is calculated according to the mean and uncertainty reported in Lu:2024ilt . The above procedure gives rise to an additional likelihood term that can be used to make inferences on the number of ER events.
For the NSI parameters, we take $\varepsilon_{\alpha\beta}$ to be uniformly distributed over the range $\left(-5,5\right)$. Moreover, we take $\varphi$ to be uniformly distributed over the range $\left(-\pi/2,\pi/2\right)$. Lastly, to ensure even sampling on the iso-$\varepsilon_{\alpha\beta}$ sphere, we set the prior on $\sin\eta$ to be distributed uniformly over $\left(-1,1\right)$. We physically motivate these priors in Appendix B. The full model is summarized below:
@@
$\begin{split}\hfill\varepsilon_{\alpha\beta}&\sim\text{Uniform}(-5,5)\,,\quad\varphi\sim\text{Uniform}(-\pi/2,\pi/2)\,,\quad\sin\eta\sim\text{Uniform}(-1,1)\,,\hfill\\ \hfill N_{\text{NR}}&\sim\text{Poisson}(\lambda_{\mathrm{bkg}}+\lambda_{\text{NR}}(\varepsilon_{\alpha\beta},\eta,\varphi))\,,\quad r_{\mathrm{ER}}\sim\text{TruncatedNormal}(1.72,1.72^{2},0,\infty)\,.\hfill\end{split}$ (8)
@@
Figure 3: The $68\%$ and $95\%$ contours for the inferred posteriors for the neutrino non-standard interactions parameters. Our sampling methods are nested sampling (NS) and the No U-Turn Sampler (NUTS), both with and without neural transport (NeuTra). Two of the six $\varepsilon_{\alpha\beta}$ are selected to be shown for visualization purposes; the full corner plot is given in Appendix E. As desired, all contours show a high degree of overlap. Probability densities are also shown with $68\%$ equal-tailed intervals indicated as dashed lines. The intervals from all methods are almost identical.
@@
Results
Method Wall Time per ESS [s] Eval. per ESS Div. per ESS [×$\mathbf{\times 10^{-4}}$]
@@ Results
NeuTra–NS $0.9\pm 0.1$ $150\pm 20$ N/A
NUTS $5.6\pm 0.2$ $195\pm 7$ $11\pm 5$
NeuTra–NUTS $2.2\pm 0.2$ $80\pm 7$ $2.3\pm 0.9$
@@
We show the marginalized posteriors for a subset of the NSI parameters in Fig. 3, produced using corner corner . As in Section 3.1, the $68\%$ and $95\%$ highest posterior density region contours, as well as the probability densities, all agree. Due to the complexity of the model, no ground truth contour is derived. The full corner plot for all NSI parameters is shown in Appendix E and all demonstrate the same level of agreement.
@@
We summarize our performances in Table 2. For the NUTS sampler, we compute the ESS separately for each dimension. We use the minimum ESS across all dimensions, which is either $\varphi$ or $\sin\eta$, to characterize the performance following hoffman2014no . We find that employing NeuTra to NUTS improves the efficiency by a factor of $2$ and reduces the number of divergences by a factor of $5$. Applying NeuTra to NS accelerates it by a larger factor of 10. We therefore find that NeuTra–NS provides the greatest performance in the NSI problem.
@@
Unlike for our synthetic problem where NeuTra–NUTS was more performant, NeuTra–NS is the better choice for our physics problem. We believe that this is due to the computational expense of gradient evaluations given the complexity of our likelihood function. This highlights how the optimal strategy for sampling statistical problems needs to be evaluated on a case-by-case basis.
@@
Nevertheless, as with the synthetic problem in Section 3.1, we find the Bayesian evidence to be consistent between NS and NeuTra-NS runs. These values are shown in Appendix F and are also close to the mean of the ELBO from the last 100 epochs of NeuTra training, $-10.04$.
@@
Lastly, our result represents the first NSI parameter space scan for direct detection experiments whereby all the parameters are allowed to vary simultaneously. A similar scan was carried out in Amaral:2023tbs ; however, only two parameters were allowed to vary at once. Our approach instead captures the full complexity of this space, allowing us to draw full multi-dimensional credible regions.
@@ 4 Conclusions
We have achieved significant performance improvements in Bayesian inference compared to techniques traditional to the astroparticle physics community. We accomplished this by leveraging GPU acceleration, automatic differentiation, and neural-network-guided reparameterization, benchmarking their performances against nested sampling alone. As our physics application, we have made inferences in the multi-dimensional parameter space of neutrino non-standard interactions using the astroparticle experiments XENON1T and PandaX-4T. We have achieved a factor $\mathord{\sim}100$ performance boost compared to nested sampling alone without compromising the accuracy of the results, also producing the first parameter space scan for direct detection experiments where all NSI parameters are allowed to vary simultaneously.
@@
Applying these methods to our astroparticle physics scenario, we found that GPUs accelerated likelihood evaluations by a factor of $\mathord{\sim}20$ and that NUTS was faster than nested sampling by a factor of $\mathord{\sim}2$. Finally, reparameterization using neural transport, whereby our posteriors were first mapped to Gaussian geometries, gave us a total factor of $\mathord{\sim}60$ improvement with NUTS and $\mathord{\sim}100$ with nested sampling compared to nested sampling on the CPU without neural transport. Unlike in the Gaussian fit, the performance of NUTS was lower than that of nested sampling when neural transport was employed, demonstrating that the optimal method depends on the specific problem at hand. In addition, since nested sampling allows for the computation of the Bayesian evidence, this work represents a way to accelerate model comparison while retaining compatibility with existing nested sampling implementations that are widely used in the natural sciences.
Our results underscore the potential of advanced computational techniques to transform inference and model comparison in astroparticle physics. The performance improvements we have achieved will aid in diagnosing convergence issues, incorporating additional experimental data, and producing new research. Moreover, these methods can be extended to other areas featuring multi-dimensional parameter spaces.
@@ 4.1 Limitations and Future Work
For our neutrino physics experiments, the computational resource used to train the neural transport model was insignificant compared to the sampling process. This may not be true if only a small posterior sample is needed, and thus whether neural transport would yield performance improvements would depend on the desired posterior sample size. We note that the effectiveness of neural transport is dependent on the ability of the trained normalizing flow to represent the desired posterior distribution pmlr-v202-grenioux23a and that Neural transport requires manual tuning for individual problems. For future work, we will assess how comparing the Bayesian evidence evaluated using nested sampling and the ELBO from neural transport training can be used as a diagnostic for both the convergence of nested sampling and the training of the normalizing flow.
Finally, we are applying the methods in this work to so-called global fits of neutrino properties. In these fits, multi-dimensional parameter spaces are independently scanned for multiple experiments, and they are produced by collaborations such as NuFIT Esteban:2020cvm . Combining the constraining power from many experiments—especially those with multiple measurement channels such as DUNE DUNE:2020ypp , PandaX PandaX:2014mem , and XENONnT XENON:2024wpa —will make our methods particularly pertinent. This is because each experiment will have an associated complex posterior geometry to traverse that can benefit from the techniques we have explored.
Acknowledgments and Disclosure of Funding
We thank Aarón Higuera for insightful discussions throughout this work. We also thank Fangda Gu, Ray Hagimoto, Aarón Higuera, Ivy Li, and Andre Scaffidi for their comments on the manuscript. This work is supported by the Department of Energy AI for HEP program, Rice University, and The National Science Foundation award PHY-204659. We thank Nvidia for supplying us with the Titan V GPUs used in this work.
@@ Appendix A Reparameterizing Probability Density Functions
$\int_{\boldsymbol{f}(\boldsymbol{\Omega})}\rho_{\boldsymbol{Y}}(\boldsymbol{y})\mathop{}\!\mathrm{d}^{n}\boldsymbol{y}=\int_{\boldsymbol{\Omega}}\rho_{\boldsymbol{Y}}[\boldsymbol{f}(\boldsymbol{x})]\left|\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|\mathop{}\!\mathrm{d}^{n}\boldsymbol{x}\,,$ (10)
@@
Appendix B Neutrino Non-Standard Interactions
Neutrino non-standard interactions (NSI) is a physics framework describing potential novel phenomena that can occur with fundamental particles known as neutrinos. We can model these phenomena via the parameterization described in Section 2.2 and visualized in Fig. 1. The physics underlying neutrino NSI can also be represented as per Fig. 4, where a neutrino of flavor $\alpha$ interacts with a particle (proton, neutron, or electron) and leaves as a neutrino of flavor $\beta$. The strength of this interaction is captured by the parameter $\varepsilon_{\alpha\beta}$, and how much of it takes place with these three particles is described by the angles $\eta$ and $\varphi$.
where the factor of $5^{1/2}$ is a convention in the field [28, 9]. Except for this factor, this system of equations can be seen to be the projection from the ‘spherical’ NSI parameterization onto a ‘Cartesian’ one. Ultimately, the predictions made by the SNuDD codebase are based on the values of these latter parameters [9, 64].
We rely on the physical intuition behind these variables when considering the priors to place on the parameters $(\varepsilon_{\alpha\beta},\eta,\varphi)$, described in Section 3.2 and given explicitly in Eq. 8. Firstly, we expect that the intrinsic strength of the new neutrino interaction for any flavor combination, captured by $\varepsilon_{\alpha\beta}$, should not have any preferred value. Hence, it should be uniformly distributed over a wide range. Secondly, given a value for this interaction strength, we do not expect any preference for the specific particle with which the neutrino interacts. Thus, we posit that the surface of the sphere in the spherical NSI space defined by the fixed radius $\varepsilon_{\alpha\beta}$ should be evenly sampled. To achieve this, we can inspect Eq. 13 to notice that these transformations are nothing but the usual spherical–Cartesian transformations with the polar angle $\theta$ replaced by the latitude angle $\eta\equiv\pi/2-\theta$. In spherical polars, we evenly populate the surface of a sphere by sampling $\cos\theta\sim\mathrm{Uniform}(-1,1)$; in our case, it thus suffices to make the replacement $\cos\theta\rightarrow\sin\eta$. Lastly, the polar angle $\varphi$ is uniformly sampled in the range $(-\pi/2,\pi/2)$ instead of the usual $(0,2\pi)$ because the ‘radius’ of the sphere, given by $\varepsilon_{\alpha\beta}$, is allowed to take on negative values.
[…]
Figure 4: The neutrino non-standard interactions (NSI) picture. A neutrino ($\nu$) of flavor $\alpha$ interacts with a proton ($p$), a neutron ($n$), or an electron ($e$) and leaves as a neutrino of flavor $\beta$. The strength of this new interaction is parameterized by the variable $\varepsilon_{\alpha\beta}$.
Appendix C Ground Truth Bayesian Evidence and Contour for Synthetic Model

### v2

@@ 4.1.1 Results
To compare the performance of NUTS and NS, both with and without NeuTra, we use the metrics of wall time per effective sample size (ESS), and likelihood or gradient evaluation per ESS. For NUTS, as ESS can be different for each parameter, the minimum ESS is used, following [19]. The calculation of ESS for MCMC and NS is explained in C. We note that the ESS per evaluation metric is not directly comparable between NS and NUTS because gradient evaluations are required for NUTS and not for NS. We also include the number of divergences encountered per ESS for NUTS runs, as these indicate regions of high curvature that often cannot be reliably sampled by MCMC [45].
We show the marginalized posteriors for $\mu_{1}$ and $\mathcal{C}_{1}$ in Fig. 2, produced using corner [71]. We see that the $68\%$ and $95\%$ highest posterior density region contours, as well as the probability densities of these parameters, all show a high degree of overlap. Our results also agree with the ground truth. The full corner plot for all parameters is shown in D, and they exhibit the same behaviors.
@@
We also find that the evidence computed using nested sampling, shown in E, matches the ground truth value of $-11.14$. The mean ELBO from the last 100 epochs of neural transport training, $-11.21$, is also close to the ground truth value. This demonstrates that NeuTra can be applied to NS to accelerate Bayesian evidence computation and model comparison.
4.2 Application to Neutrino Non-Standard Interactions
4.2.1 NSI Model
@@
We construct our likelihood using the NR result from XENON1T [34] and the ER measurement from PandaX-4T [35]. These measurements probe complementary parts of the NSI parameter space, allowing us to make stronger inferences compared to using either result alone.
@@
For the XENON1T NR measurement, the data contains both solar neutrino events and background events. As our likelihood model, we use a Poisson distribution with expectation value $\lambda\equiv\lambda_{\mathrm{NR}}+\lambda_{\mathrm{bkg}}$. Here, $\lambda_{\text{NR}}$ is the expected number of NR events, calculated using SNuDD via Eq. 7, and $\lambda_{\mathrm{bkg}}$ is the expected number of background events. As given in [34], we use $\lambda_{\mathrm{bkg}}=5.38$ and take the number of observed NR events to be $N_{\mathrm{NR}}=6$.
@@
For the PandaX-4T ER result, we use their inferred result on the number of solar neutrinos arriving at the detector per second per unit area, which is proportional to the number of ER events. As in the NR case, we use SNuDD to compute the number of expected ER events for a given set of NSI parameters. We then compare this number to the one calculated in the absence of novel phenomena using the ratio between the two values, $r_{\mathrm{ER}}$. We model the distribution of this ratio as a normal distribution truncated at zero at the lower end with mean and standard deviation both given by $1.72$. The truncation is to indicate that this physical quantity is strictly positive. The value $1.72$ is calculated according to the mean and uncertainty reported in [35]. The above procedure gives rise to an additional likelihood term that can be used to make inferences on the number of ER events.
We physically motivate the priors on the NSI parameters below. Firstly, we expect that the intrinsic strength of the new neutrino interaction for any flavor combination, captured by $\varepsilon_{\alpha\beta}$, should not have any preferred value. Hence, it should be uniformly distributed over a wide range, which we take to be $\left(-5,5\right)$. Secondly, given a value for this interaction strength, we do not expect any preference for the specific particle with which the neutrino interacts. Thus, we consider the prior to be uniformly distributed over the surface of the sphere in the spherical NSI space defined by the fixed radius $\varepsilon_{\alpha\beta}$. As such, we take $\sin\eta$ to be uniformly distributed over $(-1,1)$. The polar angle $\varphi$ is uniformly distributed in the range $(-\pi/2,\pi/2)$ instead of the usual $(0,2\pi)$ since the ‘radius’ of the sphere, given by $\varepsilon_{\alpha\beta}$, is allowed to take on negative values. The full model is summarized as follows:
@@
$\begin{split}\hfill\varepsilon_{\alpha\beta}&\sim\text{Uniform}(-5,5)\,,\\ \varphi&\sim\text{Uniform}(-\pi/2,\pi/2)\,,\\ \sin\eta&\sim\text{Uniform}(-1,1)\,,\hfill\\ \hfill N_{\text{NR}}&\sim\text{Poisson}(\lambda_{\mathrm{bkg}}+\lambda_{\text{NR}}(\varepsilon_{\alpha\beta},\eta,\varphi))\,,\\ r_{\mathrm{ER}}&\sim\text{TruncatedNormal}(1.72,1.72^{2},0,\infty)\,.\hfill\end{split}$ (9)
@@
Figure 3: The $68\%$ and $95\%$ contours for the inferred posteriors for the neutrino non-standard interactions parameters. Our sampling methods are nested sampling (NS) and the No U-Turn Sampler (NUTS), both with and without neural transport (NeuTra). Two of the six $\varepsilon_{\alpha\beta}$ are selected to be shown for visualization purposes; the full corner plot is given in D. As desired, all contours show a high degree of overlap. Probability densities are also shown with $68\%$ equal-tailed intervals indicated as dashed lines. The intervals from all methods are almost identical.
@@
4.2.2 Results
Method Wall Time/ESS [s] Eval./ESS Div./ESS [×$\mathbf{\times 10^{-4}}$]
@@ 4.2.2 Results
NeuTra–NS (GPU) $0.9\pm 0.1$ $150\pm 20$ N/A
NUTS (GPU) $5.6\pm 0.2$ $195\pm 7$ $11\pm 5$
NeuTra–NUTS (GPU) $2.2\pm 0.2$ $80\pm 7$ $2.3\pm 0.9$
@@
We show the marginalized posteriors for a subset of the NSI parameters in Fig. 3, produced using corner [71]. As in Section 4.1, the $68\%$ and $95\%$ highest posterior density region contours, as well as the probability densities, all agree. Due to the complexity of the model, no ground truth contour is derived. The full corner plot for all NSI parameters is shown in D and all demonstrate the same level of agreement.
@@
We summarize our performances in Table 2. For the NUTS sampler, we compute the ESS separately for each dimension. We use the minimum ESS across all dimensions, which is either $\varphi$ or $\sin\eta$, to characterize the performance following [19]. We find that employing NeuTra to NUTS improves the efficiency by a factor of $2$ and reduces the number of divergences by a factor of $5$. Applying NeuTra to NS accelerates it by a larger factor of 10. We therefore find that NeuTra–NS provides the greatest performance in the NSI problem.
@@
Unlike for our synthetic problem where NeuTra–NUTS was more performant, NeuTra–NS is the better choice for our NSI problem. We believe that this is partially due to the computational expense of gradient evaluations given the complexity of our likelihood function. With information about the relative cost of gradient evaluations versus likelihood calls without automatic differentiation, it is possible to use the evaluations per effective sample information in Tables 1 and 2 to estimate which method would have a shorter runtime. It should however be noted that the ratio of evaluations per effective sample does not stay constant between the two problems, indicating that the specific structure of the posterior plays a role as well. This highlights how the optimal strategy for sampling statistical problems needs to be evaluated on a case-by-case basis.
@@
Nevertheless, as with the synthetic problem in Section 4.1, we find the Bayesian evidence to be consistent between NS and NeuTra-NS runs. These values are shown in E and are also close to the mean of the ELBO from the last 100 epochs of NeuTra training, $-10.04$.
@@
Our result represents the first NSI parameter space scan for direct detection experiments whereby all the parameters are allowed to vary simultaneously. A similar scan was carried out in [9]; however, only one parameter was permitted to vary at once. Our approach instead captures the full complexity of this space, allowing us to draw full multi-dimensional credible regions. In doing so, we have taken an ‘agnostic’ EFT approach to the problem, in which we are neutral about the operators coming into play at the new physics scale, with no single one dominating. Similar studies were conducted in, for instance, Refs. [62, 72, 28]. We present the $1$D $90\%$ credible level intervals on each of our NSI parameters in Table 3, obtaining these by marginalizing over all other parameters. We use the marginalized posteriors from the result of the NeuTra-NUTS method, fully visualized in Fig. 6.
NSI Parameter Credible Interval
$\varepsilon_{ee}$ $[-3.67,3.73]$
$\varepsilon_{e\mu}$ $[-3.41,3.44]$
$\varepsilon_{e\tau}$ $[-3.33,3.40]$
$\varepsilon_{\mu\mu}$ $[-3.86,3.83]$
$\varepsilon_{\mu\tau}$ $[-3.70,3.75]$
$\varepsilon_{\tau\tau}$ $[-3.84,3.74]$
$\eta$ $[-37.6^{\circ},\,-32.0^{\circ}]$
[…]
Table 3: One-dimensional $90\%$ credible level intervals on each of our NSI parameters. The interval on any one parameter is computed following marginalization over all other parameters.
When taking our agnostic EFT approach, we find the expected result that the overall strength of neutrino NSI, characterized by the parameters $\varepsilon_{\alpha\beta}$, is poorly constrained. This is due to the large number of interferences that can occur between these parameters, leading to signal cancellations. Nevertheless, in this scheme, we are able to tightly constrain the ‘direction’ of NSI, finding narrow intervals for both $\eta$ and $\varphi$.
This latter result indicates that, within this agnostic scheme, neutrino NSI cannot take place purely with the electron ($\varphi=\pm 90^{\circ})$, nor can they take place only with up- or down-quarks (characterized at $\varphi=0^{\circ}$ by the angles $\eta\approx 26.6^{\circ}$ and $\eta\approx 63.4^{\circ}$, respectively). Instead, we are constrained to lie tightly along the proton direction in the charged plane and also along where there is a cancellation between proton and neutron NSI in the neutrino-nucleus cross section for a xenon target, occurring at $\eta\approx-35^{\circ}$ (see Ref. [9] for details). A similar result was found in Ref. [28], wherein combining data from processes sensitive to nuclear and electronic processes—in their case through results respectively from spallation source and solar neutrino/reactor experiments—led to tight constraints on the direction of NSI centred around our intervals.
However, a key difference between our results and those of Ref. [28] is that we are able to achieve these intervals using only a single type of experiment. Direct detection experiments are uniquely placed to probe both nuclear and electronic signals, allowing us to simultaneously test neutrino NSI with nucleons and electrons. While current experiments are unable to single-handedly provide leading constraints in this the NSI space, we expect that future experimental data will be an invaluable addition to global analyses [9]. In this case, having robust methods to include direct detection experiments in these studies promptly, such as those we provide in this work, will be welcomed by the community.
@@ 5 Conclusions
We have achieved significant performance improvements in Bayesian inference compared to techniques traditional to the astroparticle physics community. We accomplished this by leveraging GPU acceleration, automatic differentiation, and neural-network-guided reparameterization, benchmarking their performances against nested sampling alone. We have made inferences in the multi-dimensional parameter space of neutrino non-standard interactions using the astroparticle experiments XENON1T and PandaX-4T. We have achieved a factor $\mathord{\sim}100$ performance boost compared to nested sampling alone without compromising the accuracy of the results, also producing the first parameter space scan for direct detection experiments where all NSI parameters are allowed to vary simultaneously.
@@
Applying these methods to the NSI parameter space, we found that GPUs accelerated likelihood evaluations by a factor of $\mathord{\sim}20$ and that NUTS was faster than nested sampling by a factor of $\mathord{\sim}2$. Finally, reparameterization using neural transport, whereby our posteriors were first mapped to Gaussian geometries, gave us a total factor of $\mathord{\sim}60$ improvement with NUTS and $\mathord{\sim}100$ with nested sampling compared to nested sampling on the CPU without neural transport. Unlike in the Gaussian fit, the performance of NUTS was lower than that of nested sampling when neural transport was employed, demonstrating that the optimal method depends on the specific problem at hand. In addition, since nested sampling allows for the computation of the Bayesian evidence, this work represents a way to accelerate model comparison while retaining compatibility with existing nested sampling implementations that are widely used in the natural sciences.
Within the NSI space, this result is the first scan for direct detection experiments wherein all parameters are allowed to vary at once. We took an agnostic effective field theory approach in which not only the strength of potential new interactions with neutrinos can change but also the specific particles they can interact with can vary. We found that any individual parameter controlling the interaction strength is only loosely constrained with direct detection data, but that the particles with which new interactions can take place are well constrained at a level comparable to that achieved using dedicated neutrino experiments.
Our results underscore the potential of advanced computational techniques to transform inference and model comparison in astroparticle physics. The performance improvements we have achieved will aid in diagnosing convergence issues, incorporating additional experimental data, and producing timely studies. Moreover, these methods can be extended to other areas in astroparticle physics featuring multi-dimensional parameter spaces.
@@ 5.1 Limitations and Future Work
For our neutrino physics experiments, the computational resource used to train the neural transport model was insignificant compared to the sampling process. This may not be true if only a small posterior sample is needed, and thus whether neural transport would yield performance improvements would depend on the desired posterior sample size. We note that the effectiveness of neural transport is dependent on the ability of the trained normalizing flow to represent the desired posterior distribution [73] and that Neural transport requires manual tuning for individual problems. For future work, we will assess how comparing the Bayesian evidence evaluated using nested sampling and the ELBO from neural transport training can be used as a diagnostic for both the convergence of nested sampling and the training of the normalizing flow. In this work, while each likelihood evaluation is internally vectorized, we have also not worked on vectorizing our likelihood functions for batched evaluation. Vectorized likelihood functions might lead to GPUs having an even greater speedup over CPUs for nested sampling, and we hope to address this in the future.
Finally, we are applying the methods in this work to so-called global fits of neutrino properties. In these fits, multi-dimensional parameter spaces are independently scanned for multiple experiments, and they are produced by collaborations such as NuFIT [74]. Combining the constraining power from many experiments—especially those with multiple measurement channels such as DUNE [75], PandaX [76], and XENONnT [77]—will make our methods particularly pertinent. This is because each experiment will have an associated complex posterior geometry to traverse that can benefit from the techniques we have explored.
6 Data Availability Statement
[…]
We thank Aarón Higuera for insightful discussions throughout this work. We also thank Fangda Gu, Ray Hagimoto, Aarón Higuera, Ivy Li, and Andre Scaffidi for their comments on the manuscript. This work is supported by the Department of Energy AI4HEP program, Rice University, and the National Science Foundation CAREER award PHY-204659. We thank Nvidia for supplying us with the Titan V GPUs used in this work.
@@ Appendix A Reparameterizing Probability Density Functions
$\int_{\boldsymbol{f}(\boldsymbol{\Omega})}\rho_{\boldsymbol{Y}}(\boldsymbol{y})\,\mathrm{d}^{n}\boldsymbol{y}=\int_{\boldsymbol{\Omega}}\rho_{\boldsymbol{Y}}[\boldsymbol{f}(\boldsymbol{x})]\left|\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|\mathrm{d}^{n}\boldsymbol{x}\,,$ (11)
@@
Appendix B Ground Truth Bayesian Evidence and Contour for Synthetic Model